# Revisiting Static Feature-Based Android Malware Detection

### v1

@@ 6.3 Duplicates & Model Selection
6.4 Effective Model Selection for Robust Baselines
Our experiments with offline learning show that the XGBoost model achieves superior performance across various experimental settings, contrasting with previous results [15]. The key reason is the expanded hyperparameter search space. We included more hyperparameters and broadened the search range, as detailed in Appendix B. A critical change was the number of boosting rounds considered. Chen et al. [15] tested rounds ranging from 10 to 100. However, as shown in Figure 8, XGBoost often benefits from more boosting rounds, particularly on the APIGraph dataset, where 400 rounds yielded the best results for duplicated and deduplicated datasets. This underscores the importance of proper hyperparameter tuning to establish stronger baseline models [6].
@@ 6.4 Effective Model Selection for Robust Baselines
Table 4: Performance of different models on Drebin Dataset for active learning with 50 & 100 annotation budget per month
@@
Table 5: Performance of different models on APIGraph Dataset for active learning with 50 & 100 annotation budget per month
7 Active Learning Results & Analysis
In the active learning setting, duplicates pose additional challenges beyond those in offline learning. Since active learning relies on human experts to select samples for annotation, increasing the diversity of selected samples is crucial [48]. Duplicates can lead to redundant selections, wasting the annotation budget. A simple heuristic can identify previously annotated samples and replace them with their existing annotations. Thus, using duplicated datasets in an active learning setting is impractical, and our analysis primarily focuses on deduplicated datasets.
7.1 Malware Detection Performance
We present the results of the six models in Tables 4 and 5 for the Drebin and APIGraph datasets, discussing merged and holdout training settings. Unlike offline learning, in active learning, the model is continuously updated during validation months using a subset of samples. Key findings are summarized below:
@@ 7.1 Malware Detection Performance
    XGBoost’s Superior Performance: Similar to offline learning, XGBoost achieves the highest average F1 score across all experimental settings in active learning. This is especially evident in the challenging Drebin dataset, where it surpasses the second-best model by 8.1% in the merged training setting. This highlights the importance of selecting a strong baseline model and underscores XGBoost’s effectiveness in Android malware detection.
@@
    Performance Gains from Merged Training: Merging the validation set with the training set can enhance performance on test data, particularly for neural networks. However, the performance difference is less pronounced since models are retrained during validation months with a subset of samples.
7.2 Variance in Results
Figure 9 presents the monthly F1 scores on the Drebin dataset for active learning settings, focusing on four models: SVM, XGBoost, MLP, and HCC, in both duplicated and deduplicated datasets. Separate hyperparameter tuning was conducted for SVM, XGBoost, and MLP, while for HCC, we used the artifacts provided in [15] for the duplicated setting, as it aligns with their study.
In the deduplicated dataset, XGBoost performs best in most test months. With an F1 score difference of at least 1% considered "significant," XGBoost outperforms HCC in 15 out of 18 test months, while HCC leads in 2 months, with one month showing less than a 1% difference. On the original duplicated dataset, HCC surpasses XGBoost in the global average F1 metric by 3.4%. However, a monthly analysis reveals a different picture: XGBoost outperforms HCC in 9 months, while HCC leads in 8 months. This highlights the importance of detailed monthly analyses in comparing malware detection models rather than relying solely on a single metric [35].
8 Number of Trials and Performance Convergence
We plot the performance on the validation set as the number of hyperparameter search trials increases across different datasets. This is shown for the offline learning setting in Figure 10 and for the active learning setting in Figure 11. Most models converge by 100 trials in the offline learning setting. In the active learning setting, convergence occurs even earlier, suggesting that initial hyperparameters have less impact on the results. While the performance trend on the validation set can be a good indicator of model performance, this is not always the case. For example, the RF model seems to perform much better on the validation set of the Drebin dataset for both offline and active learning settings than on the test set. Therefore, using a separate validation split for model selection in malware detection is crucial.
@@ 8 Number of Trials and Performance Convergence
Figure 11: Performance on validation set as number of hyperparameter trials is increased for active learning
9 Recommendations for Better Reproducibility
Based on our reproducibility study of Android malware detection, we propose a set of recommendations to enhance the reproducibility of reported results. While demonstrated in the context of Android malware detection, these guidelines apply to other security domains where temporal dataset evaluation is essential for fair assessment of model performance in real-world settings. These recommendations should be used alongside existing guidelines for machine learning in security [6, 41] and Android malware detection [42] to improve reproducibility in future work.
  • Datasets should be deduplicated to provide a more accurate estimation of model performance and improve reproducubility of research.
  • Datasets should be split into non-overlapping train, validation, and test sets for model selection and hyperparameter tuning. Once tuning is complete, the validation set can be merged with the training data in offline learning settings to estimate model performance on test data.
  • Proper hyperparameter tuning for baseline machine learning models is crucial to understanding improvements introduced by novel approaches. For Android malware detection, XGBoost is a strong benchmark for offline and active learning settings.
  • For neural networks, reporting performance using multiple random seeds on the test set is important to account for variance introduced by random weight initialization.
  • Instead of relying on a single metric, a thorough analysis should be performed on results obtained on a month-by-month basis to compare different models and identify any potential bias leading to flawed conclusion.
10 Limitations
In this section, we discuss the potential limitations of our work. We focus solely on exact duplicates in model input, ignoring cases of applications with similar functionalities that may result in near duplicates in the feature space [4]. Even with this strict definition, up to 50% duplicates can be present in publicly available Android malware detection datasets. Extending the duplicate analysis to other domains, such as Windows malware, might require near-duplicate analysis, such as Windows malware [5, 54], which use different features.
Our study focused on two datasets with different feature sets, but many other feature sets and datasets have been proposed in the literature [36]. Our goal was to highlight how duplicates and other methodological pitfalls can affect results rather than providing a thorough comparison of different data and feature sets. Our findings on model selection, robust baselines, and evaluation strategies to enhance reproducibility apply to other feature sets and domains and can be applied in the future.
Datasets with labels aggregated from sources such as VirusTotal may have issues with ground truth annotations, as noted in prior works [3], but this is beyond our study’s scope. We focus on offline and active learning for malware detection. In the active learning setting, we retrain the model monthly with new samples, as done in prior work. However, this may be inefficient if model performance has not degraded, and a drift detection model may be needed to determine when retraining is necessary [30]. Our analysis can be extended to datasets and models used for such tasks.
@@ 11 Conclusion
In this paper, we have critically examined the challenges of reproducibility and replicability in Android malware detection, focusing on key pitfalls in dataset curation, model selection, and evaluation strategies. Our extensive evaluation across multiple machine learning models in offline and active learning settings reveals that issues such as data duplication, inadequate hyperparameter tuning, and biased model selection can lead to significant performance disparities. These findings highlight the need for more rigorous approaches to ensure the reliability of reported results. Our results indicate that, with proper tuning, simpler models can often match or surpass the performance of more complex methods, challenging the trend of relying solely on sophisticated architectures. To support future research, we provide detailed recommendations to enhance reproducibility in Android malware detection using ML, many of which can also be applied to other security tasks. Our open-source code offers a standardized, extensible framework for benchmarking new models, facilitating fairer and more reliable comparisons across studies. By making our framework available, we aim to contribute to a more transparent and robust research environment, thereby enhancing the reproducibility of future work in this field.
Appendix A Implementation Details
A.1 Environment
We conducted all experiments using Python 3.11.6 on Red Hat Enterprise Linux (Version 9.4). Neural network training was performed using a single NVIDIA A100 GPU. Table 6 lists the Python pip packages used in the implementation:
Pip Package Version Utility
scikit-learn 1.5.0 Used to train RF, SVM
torch 2.1.0 Used to train neural networks (MLP, SCC, HCC)
xgboost 2.1.0 Used to train XGBoost
numpy 1.26.4 Used to process datasets
pytorch-metric-learning 2.6.0 Used for creating data sampler for HCC
yacs 0.1.8 Configuration management for experiments
Table 6: Python pip packages used in the implementation
A.2 Code Structure
The code is organized into multiple key modules to facilitate easy extension:
  1. 1.
    Dataset: We preprocess data into a single .npz file containing features, labels, timestamps, and duplication indicators. The intra_split_dupes field marks duplicate within the current timestamp, referring to earlier indices or set to -1 if unique. The cross_split_dupes field tracks duplicates across splits, indicating the relevant split and index or none if unique. This format supports active and offline learning deduplication. The MalwareDataset class handles dataset loading and includes the HalfSampler method used in HCC [15], as well as a class for working with Triplet Datasets.
  2. 2.
    Models: This module implements six models, all inheriting from a BaseModel class, which takes params for hyperparameters and cfg for experiment-specific settings (e.g., output classes, GPU usage). Each model includes methods for fitting (fit), prediction (predict, predict_proba), and active learning sample selection (sample_active_learning).
  3. 3.
    Tasks: We separate offline and active learning into distinct tasks within the tasks module, each with specific training and evaluation procedures. Tasks use a yacs configuration [23], provided as a YAML file, specifying the task type, model, and relevant settings. This separation allows easy extension for other malware analysis tasks, such as family classification or concept drift detection, with minimal code changes.
Appendix B Hyperparameter Search
B.1 Hyperparameter Search-Space
The detailed hyperparameter search space is shown in Table 7. Most hyperparameter names follow standard terminologies in the scikit-learn, XGBoost, or PyTorch APIs. Below, we describe those with different or new terminology:
  • XGBoost/MLP/SCC - class-weight: This parameter, "scale_pos_weight" in XGBoost, defaults to 1 if set to False. When True, it adjusts for class imbalance by weighting the positive class (malware) according to the benign-to-malware ratio in the training data. For MLP and SCC, samples are either balanced in each batch with equal benign and malware samples (balance=True) or drawn randomly (balance=False).
  • MLP/SCC/HCC - cont_learning_epochs: In active learning with neural networks, initial and retraining epochs differ, similar to [15]. For HCC, it specifies the number of epochs; for MLP and SCC, it represents the fraction of the original training epochs.
@@ B.1 Hyperparameter Search-Space
While we mostly follow the hyperparameters for HCC from [15], there are a few minor differences. These choices ensure consistency with other neural network architectures and ease of implementation. We explore two architectures for the encoder and MLP layers, each instead of the single one used in their work. We include dropout as a hyperparameter instead of setting it to a fixed value of 0.2 in their study. Only the Adam optimizer is used during the active learning retraining phase. We employ a step learning rate scheduler and omit the cosine annealing scheduler. Regardless, we verify that our search space includes the best hyperparameter set reported in [15]; for instance, cosine annealing did not outperform step-based learning rates in any dataset in their results.
@@
class_weight {None, "balanced"}
@@
class_weight {None, "balanced"}
@@
Table 7: Hyperparameter Search Spaces for Different Models. U indicates sampling from a uniform random distribution over values in the range. * indicates hyperparameters that are only used during active learning retraining phase.
B.2 Influence of Individual Hyperparameters
Figure 12 illustrates the impact of individual hyperparameters on the validation set performance for the XGBoost model in the offline learning setting.
Figure 12: Influence of individual hyperparameters on different datasets for XGBoost

### v2

@@ 6.3 Effective Model Selection for Robust Baselines
TABLE IV: Performance of different models on the Drebin Dataset for continuous active learning with 50 & 100 annotation budget per month
@@
TABLE V: Performance of different models on the APIGraph Dataset for continuous active learning with 50 & 100 annotation budget per month
@@
6.4 Duplicates & Model Selection
The presence of duplicates can affect hyperparameter tuning and model selection. As shown in Figure 7, adjusting the scale_pos_weight parameter to balance classes improves validation performance for duplicated datasets but can reduce it on the deduplicated APIGraph dataset, indicating that duplicates influence optimal hyperparameter behavior.
Figure 8 further shows that duplicates alter feature weighting in XGBoost. Comparing models trained on deduplicated and original Drebin datasets, the top 50 features from the deduplicated model display markedly different importance distributions when evaluated on the duplicated data (feature names in Appendix A.3).
7 Continuous Active Learning Results & Analysis
In the continuous active learning setting, duplicates pose additional challenges beyond those in offline learning. Since active learning relies on human experts to select samples for annotation, increasing the diversity of selected samples is crucial [56]. Duplicates can lead to redundant selections, wasting the annotation budget. A simple heuristic can identify previously annotated samples and replace them with their existing annotations. Thus, using duplicated datasets in an active learning setting is impractical, and our analysis primarily focuses on deduplicated datasets.
We present the results of the six models in Tables IV and V for the Drebin and APIGraph datasets, discussing merged and holdout training settings. Key findings are summarized below:
@@ 7 Continuous Active Learning Results & Analysis
    XGBoost as a strong baseline in continuous active learning: XGBoost achieves the highest average F1 score across all experimental settings in continuous active learning. This is especially evident in the challenging Drebin dataset, where it surpasses the second-best model by 8.7% in the merged training setting. This highlights the importance of selecting a strong baseline model and underscores XGBoost’s effectiveness in Android malware detection.
@@
    Performance Gains from Merged Training: Merging the validation set with the training set can enhance performance on test data, particularly for neural networks. However, the performance difference is less pronounced than offline learning since models are retrained during validation months with a subset of samples.
Figure 9: Evolution of normalized XGBoost feature importance from July 2013 to June 2014.
Figure 10: Evolution of feature importance for MLP models from July 2013 to June 2014.
Figure 11: Evolution of normalized feature importance for HCC models from July 2013 to June 2014.
8 Post-hoc explainability
To examine how learned features change over time, we analyzed XGBoost, MLP, and HCC classifiers on the APIGraph dataset using SHAP-based explanations [57]. For each model, we constructed a reference set by taking the union of the top-10 most important features identified in the first and last months of the initial one-year period (July 2013–June 2014), using TreeSHAP for XGBoost and KernelSHAP for MLP and HCC. We then fixed this set as a basis for comparison and, for each subsequent month, recomputed feature importances using correctly classified malware samples from that month’s test data. All importances were normalized to ensure comparability. The resulting heatmaps (Figures 9, 10, and 11) reveal the temporal dynamics of feature relevance across models.
Figure 9 shows that in XGBoost models several features (e.g., Feature 12 [GetDeviceID], Feature 16 [GetSubscriberID], and Feature 87 [SendSMS]) remain consistently influential across months. In contrast, Feature 4 [ReadPhoneStateUsed] is prominent in the early months but declines in importance toward the end of the year, where Feature 44 [ReadPhoneStateRequested] gains greater weight.
For neural network models, the importance distribution is more static, with less variation across months. Despite their different training strategies (HCC being trained with hierarchical contrastive features), MLP and HCC show substantial overlap in their top features (e.g., the top four features are identical in the first month). While MLP consistently assigns high importance to Feature 16 [GetSubscriberID], HCC emphasizes both Feature 16 [GetSubscriberID] and Feature 87 [SendSMS] as particularly influential.
Figure 12: Comparison of attack success rate
9 Adversarial Robustness
We evaluate a pure black-box, score-based, feature-space evasion [58] on APIGraph and Drebin datasets. The Constrained Adaptive Attack (CAA) is a state-of-the-art method for adversarial evasion on tabular data that combines a fast gradient-based phase (CAPGD) with a stronger search phase. In our setting, we disable CAPGD and use only the search phase: a best-first beam search over single-bit add-only flips that batch-queries model scores, retains the lowest-score successors in a fixed-width beam (width 16; node cap 64), and halts upon reaching the benign threshold, exhausting the flip budget, or hitting the node cap. The attacker may flip up to $B$ binary features per sample under an add-only constraint; all features are treated as mutable. We attack 500 randomly sampled malware test points per model, while benign samples remain unchanged.
Results are shown in Figure 12 for budgets $B\in\{1,2,3,4,5\}$. Even with small budgets, the attack achieves high success across all models. On APIGraph, XGBoost and HCC are more robust than MLP, whereas on Drebin, XGBoost shows higher ASR than MLP for larger budgets ($B=3,4,5$), indicating that higher clean accuracy does not necessarily imply stronger adversarial robustness.
10 Limitations & Discussions
Figure 13: Inference time (in seconds) of different machine learning models on the Drebin (left) and APIGraph (right).
Computation Time. Figure 13 illustrates a comparison of the aggregate inference times for various machine learning models evaluated on the Drebin and APIGraph test datasets. The results show that neural models consistently incur higher inference latency compared to traditional models such as Random Forest (RF) and Support Vector Machine (SVM). The experiments were conducted using Python 3.11.6 on Red Hat Enterprise Linux 9.4, running on an Intel Xeon Gold CPU at 2.20 GHz.
Number of Trials and Performance Convergence. We use random search for hyperparameter tuning, following standard practice in domain generalization where training and test sets are not i.i.d. [42, 59]. This approach ensures fair comparison without biasing toward models with larger search spaces. To maintain rigor, we allocate sufficient trials for stable validation results. As shown in Figures 14 and 15, most models converge within 100 trials for offline learning and even fewer in continuous learning, indicating limited sensitivity to initial hyperparameters. While validation performance generally reflects test performance, exceptions exist, such as the RF model on the Drebin dataset.
@@ 10 Limitations & Discussions
Figure 15: Validation performance versus number of hyperparameter trials in continuous active learning.
Annotation Interval. We use a fixed interval for continuous active learning experiments, as widely adopted in prior work on Android malware detection [25, 24, 26, 35, 28, 9]. In practical deployments, concept drift detection could decide when to retrain the model [24], but this complicates comparisons across active learning methods since different models may trigger retraining at different times. Moreover, altering the retraining interval while keeping the annotation budget fixed can impact results; for example, increasing the interval to two months may reduce performance because the model is evaluated on a longer horizon of unseen data before retraining, compared to the current one-month setup.
Annotation Budget in the Presence of Duplicates. In our active learning setup, we assume a fixed annotation budget. For practical deployment, however, duplicates introduce an additional challenge: the annotation process should prioritize uncertain samples while ensuring that no sample is annotated more than once, as repeated annotations waste valuable resources. While this constraint is relatively easy to enforce for the static-analysis features considered in our study, it can be far more challenging when dealing with dynamic features [60, 17].
Duplicate detection. We treat duplicates as samples that are indistinguishable in the model’s feature space. This does not guarantee duplication in the raw input, since obfuscation, packing, and feature drift can map distinct binaries to identical features [61, 62, 63]. Accurate malware deduplication is difficult and outside our scope; our aim is to assess reproducibility under fixed feature sets and models. Because duplicate handling is induced by the feature extractor rather than the learner, fair model comparison requires deduplication on a common feature set or, at a minimum, explicit reporting of the resulting performance variance.
Label Noise and Obfuscation. Datasets with labels aggregated from sources such as VirusTotal may suffer from noise and inconsistencies in ground truth annotations, as noted in prior studies [23]. Similarly, malware developers often employ obfuscation techniques to evade detection. Both issues can affect model performance and bias comparisons across methods. Addressing these challenges is beyond the scope of our work; however, since our focus is on the reproducibility of machine learning models, we believe our findings remain applicable in such settings.
@@ 11 Conclusion
In this paper, we revisited key reproducibility and replicability challenges in static feature-based Android malware detection. We identified pitfalls in dataset curation, model selection, and evaluation, showing that issues such as data duplication, limited hyperparameter tuning, and biased baselines can cause major performance disparities. Through extensive experiments across models in offline and continuous active learning, we found that well-tuned simple models often match or surpass complex architectures. To promote transparency and fairness, we release an open-source framework for standardized, reproducible benchmarking of new models.
Appendix A Hyperparameter Search
A.1 Hyperparameter Search-Space
The detailed hyperparameter search space is shown in Table VII. Most hyperparameter names follow standard terminologies in the scikit-learn, XGBoost, or PyTorch APIs. Below, we describe those with different or new terminology:
  • XGBoost/MLP/SCC - class-weight: This parameter, ”scale_pos_weight” in XGBoost, defaults to 1 if set to False. When True, it adjusts for class imbalance by weighting the positive class (malware) according to the benign-to-malware ratio in the training data. For MLP and SCC, samples are either balanced in each batch with equal benign and malware samples (balance=True) or drawn randomly (balance=False).
  • MLP/SCC/HCC - cont_learning_epochs: In continuous active learning with neural networks, initial and retraining epochs differ, similar to [9]. For HCC, it specifies the number of epochs; for MLP and SCC, it represents the fraction of the original training epochs.
@@ A.1 Hyperparameter Search-Space
While we mostly follow the hyperparameters for HCC from [9], there are a few minor differences. These choices ensure consistency with other neural network architectures and ease of implementation. We explore two architectures for the encoder and MLP layers, each instead of the single one used in their work. We include dropout as a hyperparameter instead of setting it to a fixed value of 0.2 in their study. Only the Adam optimizer is used during the continuous learning retraining phase. We employ a step learning rate scheduler and omit the cosine annealing scheduler. Regardless, we verify that our search space includes the best hyperparameter set reported in [9]; for instance, cosine annealing did not outperform step-based learning rates in any dataset in their results.
TABLE VII: Hyperparameter Search Spaces for Different Models. U indicates sampling from a uniform random distribution over values in the range. * indicates hyperparameters that are only used during continuous learning retraining phase.
@@
class_weight {None, ”balanced”}
@@
class_weight {None, ”balanced”}
@@
subsample $x$, where $x\in\text{U}[0.8,1.0]$
colsample_bytree $x$, where $x\in\text{U}[0.8,1.0]$
@@
A.2 Variance in Results for Continuous Learning
Figure 16 shows the monthly F1 scores on the Drebin dataset in continuous active learning settings, comparing three neural networks on both duplicated and deduplicated datasets. Separate hyperparameter tuning was performed for MLP and SCC, while for HCC, we used the artifacts provided in [9] for the duplicated setting to maintain consistency with their study. Notably, we observe extreme variance in model performance in certain months (e.g., December 2020) in the duplicated datasets.
A.3 Feature Index Mapping
Tables VIII and IX show the feature index mappings for the important features used by different models on the APIGraph and Drebin datasets, respectively.
TABLE VIII: APIGraph API feature mapping
Feature ID Feature Name
2 usedpermissionslist_android.permission.get_tasks
4 usedpermissionslist_android.permission.read_phone_state
12 suspiciousapilist_landroid/telephony/telephonymanager.getdeviceid
14 suspiciousapilist_system/bin/su
16 suspiciousapilist_landroid/telephony/telephonymanager.getsubscriberid
17 suspiciousapilist_landroid/content/pm/packagemanager.getpackageinfo
21 urldomainlist_10.0.0.172
31 activitylist_com.waps.offerswebview
37 requestedpermissionlist_android.permission.receive_boot_completed
41 requestedpermissionlist_android.permission.get_tasks
44 requestedpermissionlist_android.permission.read_phone_state
46 suspiciousapilist_landroid/telephony/gsm/smsmanager.sendtextmessage
68 usedpermissionslist_android.permission.send_sms
75 suspiciousapilist_landroid/telephony/smsmanager.sendtextmessage
87 requestedpermissionlist_android.permission.send_sms
102 intentfilterlist_android.intent.action.package_added
107 intentfilterlist_android.intent.action.package_removed
131 requestedpermissionlist_com.android.launcher.permission.install_shortcut
133 requestedpermissionlist_com.android.browser.permission.read_history_bookmarks
150 intentfilterlist_android.intent.action.user_present
154 requestedpermissionlist_android.permission.read_contacts
157 usedpermissionslist_android.permission.get_accounts
192 requestedpermissionlist_android.permission.system_alert_window
197 requestedpermissionlist_android.permission.call_phone
TABLE IX: Drebin API feature mapping
Feature ID Feature Name
1 suspiciousapilist_android.content.res.assetmanager.open
2 suspiciousapilist_android.content.context.getassets
4 suspiciousapilist_android.app.application.oncreate
6 intentfilterlist_com.android.vending.install_referrer
7 intentfilterlist_android.intent.action.boot_completed
8 requestedpermissionlist_android.permission.read_external_storage
9 requestedpermissionlist_android.permission.receive_boot_completed
10 requestedpermissionlist_android.permission.system_alert_window
11 requestedpermissionlist_android.permission.get_tasks
12 requestedpermissionlist_android.permission.write_settings
13 requestedpermissionlist_android.permission.access_wifi_state
14 requestedpermissionlist_android.permission.read_phone_state
15 requestedpermissionlist_android.permission.change_wifi_state
16 requestedpermissionlist_android.permission.write_external_storage
17 suspiciousapilist_android.app.alertdialog\\$builder.setcancelable
18 suspiciousapilist_android.app.alertdialog\\$builder.setmessage
21 suspiciousapilist_android.view.window.settype
22 suspiciousapilist_android.content.res.resources.getassets
25 intentfilterlist_android.intent.action.view
26 requestedpermissionlist_android.permission.access_coarse_location
27 requestedpermissionlist_android.permission.camera
28 requestedpermissionlist_android.permission.access_fine_location
29 requestedpermissionlist_android.permission.vibrate
35 usedpermissionslist_android.permission.access_network_state
44 suspiciousapilist_android.app.application.getpackagename
182 suspiciousapilist_android.content.context.getsharedpreferences
904 suspiciousapilist_android.app.application.registeractivitylifecyclecallbacks
1468 suspiciousapilist_android.content.intent.setflags
1615 suspiciousapilist_android.widget.toast.show
2345 suspiciousapilist_android.content.pm.packagemanager\\$namenotfoundexception.printstacktrace
2744 suspiciousapilist_android.telephony.telephonymanager.getdeviceid
3279 suspiciousapilist_ljava/io/ioexception;->printstacktrace
3727 suspiciousapilist_android.net.wifi.wifiinfo.getmacaddress
4256 suspiciousapilist_android.content.context.getdir
4665 suspiciousapilist_android.view.layoutinflater.from
5358 suspiciousapilist_ljava/lang/runtime;->exec
6216 intentfilterlist_com.google.android.c2dm.intent.receive
6218 contentproviderlist_android.support.v4.content.fileprovider
6219 hardwarecomponentslist_android.hardware.camera.autofocus
6220 hardwarecomponentslist_android.hardware.camera
6274 suspiciousapilist_android.view.surfaceview.<init>
6402 suspiciousapilist_android.app.activitymanager.getrunningservices
6405 suspiciousapilist_android.content.pm.packagemanager.getinstalledpackages
6475 suspiciousapilist_android.content.pm.applicationinfo.loadlabel
6589 activitylist_com.google.android.gms.ads.adactivity
6591 activitylist_com.google.android.gms.common.api.googleapiactivity
6592 requestedpermissionlist_android.permission.get_accounts
12412 requestedpermissionlist_com.android.vending.billing
15566 suspiciousapilist_android.content.pm.packagemanager.setcomponentenabledsetting
16973 contentproviderlist_mono.monoruntimeprovider